Since 2000/10/29, Updated 2005/03/07
[Go back to TAMA HOME] [Go back to Detector HOME]
[DT1] [DT2] [DT3] [DT4] [DT5] [DT6] [DT7] [DT8] [DT9] (Global)
[DT1] [DT2] [DT3] [DT4] [DT5] [DT6] [DT7] [DT8] [DT9] (Internal)
>> Recent Topics <<
- 2004/01/10 DT9 has been finished!
- 2003/11/29 DT9 started! [DT9 page] [DT9 internal]
- 2003/11/04 Improvement of the sensitivity
The floor level reached at 6x10-19 m/Hz1firstname.lastname@example.orgHz.
>> Sensitivity & Stability<<
- Floor sensitivity:
6x10-19 m/Hz1/2@800kHz (=2.0x10-21 /Hz1/2 in strain) [PDF]
- History of the sensitivity [PDF]
- Observable Distance for binary inspirals [PDF]
- Stability: Longest continuous lock streatches for more than 24hours [PDF]
>> Past Topics <<
- 2003/07/12 Improvement of the sensitivity
- The improvement of the sensitivity at around 100Hz by a factor of 10 was achieved, preserving the floor level of 8x10-19 m/Hz1email@example.comHz.
- 2003/04/15 DT8 has been finished!
- Observation data for 1158hours was accumulated. This corresponds to the duty cycle of 81.3%.
- 2003/02/14 8th Data Taking run (DT8) started!
- This run was a first full-time joint observation with LIGO's three interferometer. [To DT8 page]
- 2002/08/31 Improvement of the floor level
- 1.0x10-18 m/Hz1/2@1kHz with further improved sensitivity above 1kHz. (h=3x10-21/Hz1/2)
- 2002/08/24 Improved sensitivity[PDF] [EPS]
- Improvement at around 100Hz owing to narrow band alignment filter.
- 2002/08/05 Improved sensitivity
- 1.2x10-18 m/Hz1/2@1kHz with further improved sensitivity above 1kHz. (h=4x10-21/Hz1/2)
- 2002/07/31 Improved sensitivity
- 1.5x10-18 m/Hz1/2@700Hz with improved sensitivity above 1kHz.
- 2002/07/06~08 Test Run 3 (TR2)
- Continuous lock of 4hours
- 2002/05/09~10 Test Run 2 (TR2)
- 2002/04/10 Recycled Interferometer locked with the high power laser
- The current output power of the laser is 8.6W. The power incident on the
MC is 6W.
- 2002/02/13 Improved stability and sensitivity:
- continuous lock of 2h03m [PDF]
- 4.5x10-18 m/Hz1/2[PDF]
- 2002/02/01 Improved sensitivity 1.0x10-17 m/Hz1/2
- 2002/01/23 46min. of the full lock, improved sensitivity 1.0x10-17 m/Hz1/2
- 2002/01/09 Autoalignment control engaged
- 2002/01/07 Improved stability of the full lock [EPS] [PDF]
- 2001/12/24 First full lock of recycled TAMA300 [EPS] [PDF]
- 2001/12/10 First lock of power-recycled Michelson [EPS] [PDF]
- 2001/10 Recycling1 Experiment started